# A motivic Fundamental Lemma

### v3

@@ 6.4. Back to orbital integrals
It is explained in [37, Section 8.6] how to construct from the initial data $G,H,a_{H},a,\kappa$ of Theorem 6.1.1 a curve $X$ over $K$, a line bundle $D$ on $X$, a distinguished $K$-point $\nu_{0}$ and all the auxiliary data to construct the Hitchin fibrations for $G$ and $H$. Furthermore, by choosing the degree of $D$ sufficiently large, the affine linear sub-space $\mathrm{Z}\subseteq\operatorname{\mathsf{A}}_{H}$, consisting of elements $a^{\prime}\in\operatorname{\mathsf{A}}_{H}$ such that $a^{\prime}_{\nu_{0}}\equiv a_{H}\mod(T^{N})$ for $N$ as in [37, Proposition 3.5.1], is non-empty and its pull-back $\widetilde{\mathrm{Z}}\subseteq\widetilde{\operatorname{\mathsf{A}}}_{H}$ is geometrically irreducible by the same argument as in [37, Lemma 5.3.2]. For dimension reasons also $\widetilde{\mathrm{Z}}\cap\widetilde{\operatorname{\mathsf{A}}}^{{\mathrm{ani}}}_{H}$ is non-empty and geometrically irreducible. Since $K$ is pseudo-finite we deduce $\widetilde{\mathrm{Z}}\cap\widetilde{\operatorname{\mathsf{A}}}^{{\mathrm{ani}}}_{H}(K)\neq\emptyset$. Thus we may find $a^{\prime}_{H}\in\operatorname{\mathsf{A}}_{H}^{{\mathrm{ani}}}(K)$ which admits a lift to $\widetilde{\operatorname{\mathsf{A}}}^{{\mathrm{ani}}}_{H}$ and such that
@@
One can further adjust the choice of $a^{\prime}_{H}$ as in [37, Section 8.6.6] in such a way, that for all $\nu\in X\setminus U$ different from $\nu_{0}$ the Springer fibers are of a very simple form, which is analyzed in [37, Lemme 8.5.7]. In the first two cases in loc. cit. both Springer-fibers are 00-dimensional, $r^{G}_{H}(a_{H,\nu})=0$ and
@@
In the third case we still have $\chi_{\mathrm{Psf}}({\mathbf{\Lambda}}_{\nu}\backslash{\mathcal{N}}_{a^{\prime}_{H,{\nu}}},\mathrm{stab})=1$ but $r^{G}_{H}(a^{\prime}_{H,\nu})=1$. The geometry of ${\mathcal{N}}_{a^{\prime}_{\nu}}$ is described in [37, Section 8.3] and reduced to the case when ${\mathcal{N}}_{a^{\prime}_{\nu},\overline{K}}$ is an infinite chain of ${\mathbb{P}}^{1}$’s. The Kostant section provides a $K_{\nu}$-rational point for ${\mathcal{N}}_{a^{\prime}_{\nu}}$ and hence ${\mathbf{\Lambda}}_{\nu}\backslash{\mathcal{N}}_{a^{\prime}_{\nu}}$ is isomorphic to the Weil restriction along $K_{\nu}/K$ of either a nodal ${\mathbb{P}}^{1}$ with trivial $\pi_{0}(a^{\prime}_{\nu})$-action or two ${\mathbb{P}}^{1}$’s meeting in two points and $\pi_{0}(a^{\prime}_{\nu})\cong{\mathbb{Z}}/2{\mathbb{Z}}$ exchanging components and intersection points, see [37, Example 8.1.15]. It follows that in both cases $\chi_{\mathrm{Psf}}({\mathbf{\Lambda}}_{\nu}\backslash{\mathcal{N}}_{a^{\prime}_{\nu}},\kappa)={\mathbb{L}}^{\deg(K_{\nu}/K)}$. Since $r_{H}^{G}(D)=\sum_{\nu\in X\setminus U}\deg(K_{\nu}/K)r^{G}_{H}(a^{\prime}_{H,\nu})$, the combination of (6.4.1) with (6.2.3) finishes the proof of Theorem 6.1.1.

### v4

@@ 6.4. Back to orbital integrals
It is explained in [38, Section 8.6] how to construct from the initial data $G,H,a_{H},a,\kappa$ of Theorem 6.1.1 a curve $X$ over $K$, a line bundle $D$ on $X$, a distinguished $K$-point $\nu_{0}$ and all the auxiliary data to construct the Hitchin fibrations for $G$ and $H$. Furthermore, by choosing the degree of $D$ sufficiently large, the affine linear sub-space $\mathrm{Z}\subseteq\operatorname{\mathsf{A}}_{H}$, consisting of elements $a^{\prime}\in\operatorname{\mathsf{A}}_{H}$ such that $a^{\prime}_{\nu_{0}}\equiv a_{H}\mod(T^{N})$ for $N$ as in [38, Proposition 3.5.1], is non-empty and its pull-back $\widetilde{\mathrm{Z}}\subseteq\widetilde{\operatorname{\mathsf{A}}}_{H}$ is geometrically irreducible by the same argument as in [38, Lemma 5.3.2]. For dimension reasons also $\widetilde{\mathrm{Z}}\cap\widetilde{\operatorname{\mathsf{A}}}^{{\mathrm{ani}}}_{H}$ is non-empty and geometrically irreducible. Since $K$ is pseudo-finite we deduce $\widetilde{\mathrm{Z}}\cap\widetilde{\operatorname{\mathsf{A}}}^{{\mathrm{ani}}}_{H}(K)\neq\emptyset$. Thus we may find $a^{\prime}_{H}\in\operatorname{\mathsf{A}}_{H}^{{\mathrm{ani}}}(K)$ which admits a lift to $\widetilde{\operatorname{\mathsf{A}}}^{{\mathrm{ani}}}_{H}$ and such that
@@
One can further adjust the choice of $a^{\prime}_{H}$ as in [38, Section 8.6.6] in such a way, that for all $\nu\in X\setminus U$ different from $\nu_{0}$ the Springer fibers are of a very simple form, which is analyzed in [38, Lemme 8.5.7]. In the first two cases in loc. cit. both Springer-fibers are 0-dimensional, $r^{G}_{H}(a_{H,\nu})=0$ and
@@
In the third case we still have $\chi_{\mathrm{Psf}}({\mathbf{\Lambda}}_{\nu}\backslash{\mathcal{N}}_{a^{\prime}_{H,{\nu}}},\mathrm{stab})=1$ but $r^{G}_{H}(a^{\prime}_{H,\nu})=1$. The geometry of ${\mathcal{N}}_{a^{\prime}_{\nu}}$ is described in [38, Section 8.3] and reduced to the case when ${\mathcal{N}}_{a^{\prime}_{\nu},\overline{K}}$ is an infinite chain of ${\mathbb{P}}^{1}$’s. The Kostant section provides a $K_{\nu}$-rational point for ${\mathcal{N}}_{a^{\prime}_{\nu}}$ and hence ${\mathbf{\Lambda}}_{\nu}\backslash{\mathcal{N}}_{a^{\prime}_{\nu}}$ is isomorphic to the Weil restriction along $K_{\nu}/K$ of either a nodal ${\mathbb{P}}^{1}$ with trivial $\pi_{0}(a^{\prime}_{\nu})$-action or two ${\mathbb{P}}^{1}$’s meeting in two points and $\pi_{0}(a^{\prime}_{\nu})\cong{\mathbb{Z}}/2{\mathbb{Z}}$ exchanging components and intersection points, see [38, Example 8.1.15]. It follows that in both cases $\chi_{\mathrm{Psf}}({\mathbf{\Lambda}}_{\nu}\backslash{\mathcal{N}}_{a^{\prime}_{\nu}},\kappa)={\mathbb{L}}^{\deg(K_{\nu}/K)}$. Since $r_{H}^{G}(D)=\sum_{\nu\in X\setminus U}\deg(K_{\nu}/K)r^{G}_{H}(a^{\prime}_{H,\nu})$, the combination of (6.4.1) with (6.2.3) finishes the proof of Theorem 6.1.1.